Unit 24 Homework Assignment  Physics141-Summer 2009

Unit 24 homework after session 1

• Read chapter 28 in the text book Understanding Physics.

• Work textbook problems 28-13, 28-16, 28-20. Unit 24 homework after session 2

• Read chapter 28 in the text book Understanding Physics.

• Work textbook problems 28-31, 28-49. Unit 24 homework after session 3 (We are skipping session 3

but do these problems anyway.

• Read chapter 28 in the text book Understanding Physics.

• Work textbook problems 28-46, 28-48.

All Unit 24 homework problems and the unit 24 activity guide are due on the first day of Unit 25.